# Neurological disorders leading to mechanical dysfunction of the esophagus: an emergent behavior of a neuromechanical dynamical system

### 28 Feb 2024 v1

@@ 1 Introduction
Clinically, esophageal motility has been investigated with high-resolution manometry, a technology utilizing closely spaced pressure transducers positioned along the length of the esophagus (32). High-resolution manometry is an excellent method for quantifying the strength and timing of esophageal contractions (33, 34). However, it yields minimal information on inhibition other than in tonically contracted sphincters [NEW CITATION](35, 33, 34).
@@ 2.3 Esophageal motility disorders
Repetitive retrograde contractions emerge with decreased innervation of inhibitory neurons via mechanoreceptors ($S_{I}$) or increased innervation of excitatory neurons via mechanoreceptors ($S_{E}$) (Fig. 3(i)). The retrograde pattern results from a phase shift in the excitatory signal, opposing the antegrade pattern. Importantly, this phenomenon is not due to a retrograde traveling signal, as the innervation pathways remain unidirectional. Further details on reversing propagation direction are discussed extensively in our recent work [PRL MANUSCRIPT].
@@ 3 Discussion
Advancing towards experimental validation can be achieved through additional pharmacological studies, electrophysiological recordings, and by studying the neural circuity at the molecular level though immunohistochemical, optogenetics, neuronal tracing, and calcium imaging (22, 23, 24, 70, 71, 72, 73, 74). A recent study on the peristaltic reflex of the colon provides a promising template, revealing the physiological and neural elements involved using multiple experimental methods (74).
@@
Nevertheless, despite these challenges, our work opens a new area of investigation in esophagology that is based on mechanophysiology (mechanics-based organ function). One of the leading diagnostic protocols for EMDs is the Chicago Classification scheme (34). This work is a milestone in our effort to develop the first mechanics-guided disease classification and diagnostic protocol for EMDs (75, 76).
There is also a potential for broad impact beyond the esophagus. Mechanical dysfunctions caused by neurological disorders expand beyond the gastrointestinal tract and are prominent in other organs where an emergent behavior occurs such as the heart, lungs, bladder, and uterus (77, 1, 78, 79, 80). For example, acute respiratory distress syndrome and chronic obstructive pulmonary disease are associated with disturbance in the lungs’ rhythmic function (4, 81). In the heart, asynchronous in electrical activation can cause abnormalities in perfusion and pump function, leading to arrhythmias or even heart failure (82, 83, 84).
@@ 4.1 Overview of mathematical model
The esophagus is a multilayered tubular organ that connects between the mouth and the stomach, transporting swallowed material through peristaltic contraction (58). This coordinated movement results from sequential contraction of the circular muscle, that is neurally controlled (31, 85, 86). Peristalsis can be triggered by swallowing, initiated by the central nervous system (primary peristalsis) (57, 25), or locally, provoked by esophageal distention and is independent of the central nervous system (secondary peristalsis) (22, 30). Since esophageal response to FLIP is known to be involuntary and assumed to be triggered locally, the model is such that the innervation is lead by local feedback (41, 39, 87).
@@
Excitatory input from stretch. When the FLIP bag inside the esophagus is inflated, it causes the esophagus to distend, activating the proprioceptive channel that is mediated by stretch receptors, which respond to changes in strain (89, 90). These mechanoreceptors are parts of the esophageal motor neurons and are distributed along the length of the esophagus. Thus, in our neurological model, mechanoreceptors are activated when they sense that local strain increases above a threshold and send excitatory signals to both excitatory and inhibitory neuronal populations.
@@
Input from adjacent neuronal populations. It is believed that depolarization of one muscle cell in the esophagus will result in electronic increase of neighboring muscle cells in an aboral direction (29, 91). Hence, excitatory populations in the model also receive signals from adjacent proximal populations. Nearest-neighbor connections of two types are introduced (92, 93, 94). First, unidirectional, aboral excitatory connections between excitatory populations of neighboring segments leading to a propagation of activity along the chain of segments (parameter $b$ in Fig. 1(a)). Second, inhibitory connections from the inhibitory population in one segment to the neighboring (anterior) excitatory populations (parameter $d$ in Fig. 1(a)).
@@
Our study is built upon and extends the models recently developed for larval locomotion which used the Wilson-Cowan model (92, 93, 48). It is the first to connect crawling in Drosophila models to the contraction pattern of the esophagus. However, the biggest and most unique advance is in how the initial excitation occurs and the way in which the rhythmic motion emerges. In the locomotion models, the forward waves were induced by applying an external input to distal excitatory population. In the current model, no external neuronal signaling is needed to initiate the contraction nor to sustain the rhythmic contractions.
@@ 4.2 Body mechanics and fluid equations
The flow inside the FLIP device placed in the esophageal lumen is modeled as a one-dimensional, fluid-filled, flexible tube that is closed on both ends (88, 96, 97). The conservation of mass and momentum equations are
@@
to complete the system (99, 100). Here, ${P}_{o}$, $K_{e}$, and $A_{o}$ are external pressure, tube stiffness, and undeformed reference area (cross-sectional area of the tube when $P=P_{o}$), respectively. Lastly, the neurally controlled contraction of the esophageal lumen is set through dynamically varying the rest cross-sectional area of the tube by a factor $\theta(x,t)$ (98, 101, 102). $\theta(x,t)$ captures the muscular dynamic, induced by the neural activity discussed next.
@@ 4.3 Neural circuitry: excitation, inhibition, and muscle contraction pattern
Relaxation oscillators are often used to model cardiac and gastro-intestinal electrical activity (95, 103, 104, 2). To represent the electrical activity of the entire organ, we introduce a system of locally coupled Wilson–Cowan oscillators, distributed uniformly along the length of the esophagus. Each unit consists of two neuronal populations, excitatory ($E$) and inhibitory ($I$) (Fig. 1(a)). The values of $E_{i}(t)$ and $I_{i}(t)$ at each oscillator depict the activity levels of the excitatory and inhibitory neuron populations. The differential equations for the time-dependent variation of averaged excitatory and inhibitory neuronal activities at node $i$ introduced by (48) with the unidirectional coupling introduced by (93) is
@@
where $\lambda$ is activation speed (slope of sigmoid), and $\phi$ is activation threshold (location of sigmoid’s maximum slope). Lastly, $S_{E,i}$ and $S_{I,i}$ symbolize the local excitatory inputs to each population at oscillator $i$, capturing the mechano-sensory feedback coming from the stretch receptors (92, 93), discussed next.
@@ 4.4 Non-dimensional form
where $\alpha$, $\tau$, $U$, $p$, and $\chi$ are the non-dimensional variables for cross-sectional area, time, velocity, pressure, and position, respectively (88, 96). Thus, dimensional equations ((1), (2), (3), (4), (5), and (8)) become
@@ 4.5 Initial and boundary conditions
where $S_{\text{IC}}=\text{Volume}/A_{o}$ (105, 97). Additionally,
@@ 4.6 Numerical implementation
Table 1 lists the model’s parameters and their values. Values for the parameters defining esophageal body mechanics and fluid properties are approximated based on clinical data and previous computational studies (106, 107, 108, 109). The parameters involved in the neural equations are chosen such that the system has a single unstable fixed point and a stable limit cycle in response to the constant external stimulus $S_{E}$ and $S_{I}$. Depending on the values of the parameters in the system, one can obtain multiple equilibria with different stability properties. For simplicity, we choose the values introduced by (48) with small variations. For example, theoretical work concerning Wilson-Cowan oscillators often set $S_{I}~{}=0$ for simplicity (48). By setting $S_{I}\neq 0$, we ensure that both excitatory and inhibitory populations ”turn-on” when distension is applied, independent of one other. A theoretical overview of relaxation oscillators, limit cycle solution, and the constrains introduced by modeling esophageal response is provided in Section S5.2 of the supplementary.
Note that obtaining a limit cycle solution is not the only constrain in this case. The parameters (specially $b$, $d$, $w_{E}$, and $w_{I}$) are adjusted such that the model captures the qualitative patterns observed in clinical esophageal distension studies. For instance, high value for $b$ makes the propagation more distinct and the system more robust (92, 93). However, it also breaks the constrain in which the rhythmic contractions only take place along the distended section. For high $b$, excitation from proximal cells dominate, allowing the wave to propagate even without local excitation from mechanoreceptors.
@@ Acknowledgments
Lastly, note that the system’s response to the change of strength of input parameter to excitatory population, $w_{E}$, is as the center of a different study, which we encourage the readers to visit [NEURO-TECHNICAL].
@@
A relaxation oscillator is a mathematical model of electronic circuit that can exhibit a variety of behaviors that depend on the intrinsic and input parameters defining it (110). Relaxation oscillators have been used for years to model the electrical activity of the heart due to their ability to display rhythmic, repetitive patterns known as limit cycles (1, 111, 2). The shape, frequency, and amplitude of the oscillations depend on the properties of the oscillator (110).
@@
For additional information on dynamic systems, refer to (110, 52, 112).
A chain of limit cycle oscillators is a well-studied system (113, 114). The oscillators in such a system are coupled, often identical, and each oscillator exhibits a limit cycle solution, independent of the coupling (55, 115, 53). Such systems often exhibit “frequency pulling”, where all oscillators converge into an entrained frequency. However, they are not synchronized, as there is phase shift between adjacent oscillators (113, 116). At steady state, the oscillators phase lock, meaning that they oscillate at constant phase from one another, creating a special case of synchrony (49, 117). This emerging pattern, known to be the fundamental solution of coupled system in equilibrium state, is the key element in creating the propagating attribute of the system (50, 54).

### 01 Mar 2024 v2

@@ 1 Introduction
Clinically, esophageal motility has been investigated with high-resolution manometry, a technology utilizing closely spaced pressure transducers positioned along the length of the esophagus (32). High-resolution manometry is an excellent method for quantifying the strength and timing of esophageal contractions (33, 34). However it cannot yield specific information about inhibition other than in tonically contracted sphincter (35, 33, 34).
@@ 2.3 Esophageal motility disorders
Repetitive retrograde contractions emerge with decreased innervation of inhibitory neurons via mechanoreceptors ($S_{I}$) or increased innervation of excitatory neurons via mechanoreceptors ($S_{E}$) (Fig. 3(i)). The retrograde pattern results from a phase shift in the excitatory signal, opposing the antegrade pattern. Importantly, this phenomenon is not due to a retrograde traveling signal, as the innervation pathways remain unidirectional. Further details on reversing propagation direction are discussed extensively in our recent work (69).
@@ 3 Discussion
Advancing towards experimental validation can be achieved through additional pharmacological studies, electrophysiological recordings, and by studying the neural circuity at the molecular level though immunohistochemical, optogenetics, neuronal tracing, and calcium imaging (22, 23, 24, 71, 72, 73, 74, 75). A recent study on the peristaltic reflex of the colon provides a promising template, revealing the physiological and neural elements involved using multiple experimental methods (75).
@@
Nevertheless, despite these challenges, our work opens a new area of investigation in esophagology that is based on mechanophysiology (mechanics-based organ function). One of the leading diagnostic protocols for EMDs is the Chicago Classification scheme (34). This work is a milestone in our effort to develop the first mechanics-guided disease classification and diagnostic protocol for EMDs (76, 77).
There is also a potential for broad impact beyond the esophagus. Mechanical dysfunctions caused by neurological disorders expand beyond the gastrointestinal tract and are prominent in other organs where an emergent behavior occurs such as the heart, lungs, bladder, and uterus (78, 1, 79, 80, 81). For example, acute respiratory distress syndrome and chronic obstructive pulmonary disease are associated with disturbance in the lungs’ rhythmic function (4, 82). In the heart, asynchronous in electrical activation can cause abnormalities in perfusion and pump function, leading to arrhythmias or even heart failure (83, 84, 85).
@@ 4.1 Overview of mathematical model
The esophagus is a multilayered tubular organ that connects between the mouth and the stomach, transporting swallowed material through peristaltic contraction (58). This coordinated movement results from sequential contraction of the circular muscle, that is neurally controlled (31, 86, 87). Peristalsis can be triggered by swallowing, initiated by the central nervous system (primary peristalsis) (57, 25), or locally, provoked by esophageal distention and is independent of the central nervous system (secondary peristalsis) (22, 30). Since esophageal response to FLIP is known to be involuntary and assumed to be triggered locally, the model is such that the innervation is lead by local feedback (41, 39, 88).
@@
Excitatory input from stretch. When the FLIP bag inside the esophagus is inflated, it causes the esophagus to distend, activating the proprioceptive channel that is mediated by stretch receptors, which respond to changes in strain (90, 91). These mechanoreceptors are parts of the esophageal motor neurons and are distributed along the length of the esophagus. Thus, in our neurological model, mechanoreceptors are activated when they sense that local strain increases above a threshold and send excitatory signals to both excitatory and inhibitory neuronal populations.
@@
Input from adjacent neuronal populations. It is believed that depolarization of one muscle cell in the esophagus will result in electronic increase of neighboring muscle cells in an aboral direction (29, 92). Hence, excitatory populations in the model also receive signals from adjacent proximal populations. Nearest-neighbor connections of two types are introduced (93, 94, 95). First, unidirectional, aboral excitatory connections between excitatory populations of neighboring segments leading to a propagation of activity along the chain of segments (parameter $b$ in Fig. 1(a)). Second, inhibitory connections from the inhibitory population in one segment to the neighboring (anterior) excitatory populations (parameter $d$ in Fig. 1(a)).
@@
Our study is built upon and extends the models recently developed for larval locomotion which used the Wilson-Cowan model (93, 94, 48). It is the first to connect crawling in Drosophila models to the contraction pattern of the esophagus. However, the biggest and most unique advance is in how the initial excitation occurs and the way in which the rhythmic motion emerges. In the locomotion models, the forward waves were induced by applying an external input to distal excitatory population. In the current model, no external neuronal signaling is needed to initiate the contraction nor to sustain the rhythmic contractions.
@@ 4.2 Body mechanics and fluid equations
The flow inside the FLIP device placed in the esophageal lumen is modeled as a one-dimensional, fluid-filled, flexible tube that is closed on both ends (89, 97, 98). The conservation of mass and momentum equations are
@@
to complete the system (100, 101). Here, ${P}_{o}$, $K_{e}$, and $A_{o}$ are external pressure, tube stiffness, and undeformed reference area (cross-sectional area of the tube when $P=P_{o}$), respectively. Lastly, the neurally controlled contraction of the esophageal lumen is set through dynamically varying the rest cross-sectional area of the tube by a factor $\theta(x,t)$ (99, 102, 103). $\theta(x,t)$ captures the muscular dynamic, induced by the neural activity discussed next.
@@ 4.3 Neural circuitry: excitation, inhibition, and muscle contraction pattern
Relaxation oscillators are often used to model cardiac and gastro-intestinal electrical activity (96, 104, 105, 2). To represent the electrical activity of the entire organ, we introduce a system of locally coupled Wilson–Cowan oscillators, distributed uniformly along the length of the esophagus. Each unit consists of two neuronal populations, excitatory ($E$) and inhibitory ($I$) (Fig. 1(a)). The values of $E_{i}(t)$ and $I_{i}(t)$ at each oscillator depict the activity levels of the excitatory and inhibitory neuron populations. The differential equations for the time-dependent variation of averaged excitatory and inhibitory neuronal activities at node $i$ introduced by (48) with the unidirectional coupling introduced by (94) is
@@
where $\lambda$ is activation speed (slope of sigmoid), and $\phi$ is activation threshold (location of sigmoid’s maximum slope). Lastly, $S_{E,i}$ and $S_{I,i}$ symbolize the local excitatory inputs to each population at oscillator $i$, capturing the mechano-sensory feedback coming from the stretch receptors (93, 94), discussed next.
@@ 4.4 Non-dimensional form
where $\alpha$, $\tau$, $U$, $p$, and $\chi$ are the non-dimensional variables for cross-sectional area, time, velocity, pressure, and position, respectively (89, 97). Thus, dimensional equations ((1), (2), (3), (4), (5), and (8)) become
@@ 4.5 Initial and boundary conditions
where $S_{\text{IC}}=\text{Volume}/A_{o}$ (106, 98). Additionally,
@@ 4.6 Numerical implementation
Table 1 lists the model’s parameters and their values. Values for the parameters defining esophageal body mechanics and fluid properties are approximated based on clinical data and previous computational studies (107, 108, 109, 110). The parameters involved in the neural equations are chosen such that the system has a single unstable fixed point and a stable limit cycle in response to the constant external stimulus $S_{E}$ and $S_{I}$. Depending on the values of the parameters in the system, one can obtain multiple equilibria with different stability properties. For simplicity, we choose the values introduced by (48) with small variations. For example, theoretical work concerning Wilson-Cowan oscillators often set $S_{I}~{}=0$ for simplicity (48). By setting $S_{I}\neq 0$, we ensure that both excitatory and inhibitory populations ”turn-on” when distension is applied, independent of one other. A theoretical overview of relaxation oscillators, limit cycle solution, and the constrains introduced by modeling esophageal response is provided in Section S5.2 of the supplementary.
Note that obtaining a limit cycle solution is not the only constrain in this case. The parameters (specially $b$, $d$, $w_{E}$, and $w_{I}$) are adjusted such that the model captures the qualitative patterns observed in clinical esophageal distension studies. For instance, high value for $b$ makes the propagation more distinct and the system more robust (93, 94). However, it also breaks the constrain in which the rhythmic contractions only take place along the distended section. For high $b$, excitation from proximal cells dominate, allowing the wave to propagate even without local excitation from mechanoreceptors.
@@ Acknowledgments
Lastly, note that the system’s response to the change of strength of input parameter to excitatory population, $w_{E}$, is as the center of a different study, which we encourage the readers to visit (69).
@@
A relaxation oscillator is a mathematical model of electronic circuit that can exhibit a variety of behaviors that depend on the intrinsic and input parameters defining it (111). Relaxation oscillators have been used for years to model the electrical activity of the heart due to their ability to display rhythmic, repetitive patterns known as limit cycles (1, 112, 2). The shape, frequency, and amplitude of the oscillations depend on the properties of the oscillator (111).
@@
For additional information on dynamic systems, refer to (111, 52, 113).
A chain of limit cycle oscillators is a well-studied system (114, 115). The oscillators in such a system are coupled, often identical, and each oscillator exhibits a limit cycle solution, independent of the coupling (55, 116, 53). Such systems often exhibit “frequency pulling”, where all oscillators converge into an entrained frequency. However, they are not synchronized, as there is phase shift between adjacent oscillators (114, 117). At steady state, the oscillators phase lock, meaning that they oscillate at constant phase from one another, creating a special case of synchrony (49, 118). This emerging pattern, known to be the fundamental solution of coupled system in equilibrium state, is the key element in creating the propagating attribute of the system (50, 54).